OpenStudy (anonymous):

Need some assistance if anyone is available: Problem #7: At time t in seconds, a particle's distance s(t), in micrometers (um), from a point is given by s(t)=e^(t)-1. What is the average velocity of the particle from t=2 to t=4? If anyone has some time right now, i'd appreciate it if someone could work through this problem with me.

OpenStudy (cwrw238):

ok first find s at time t = 2 seconds then at time t = 4 secs then subtract the distances s(4) - s(2) and divide by 4-2 = 2 seconds average velocity = (s(4) - s(2)) / 2

OpenStudy (anonymous):

just to clear things up, to find s at time t=2 i plug t=2 into s(t)=x^(t)-1 correct?

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

alright let me try all that out.

OpenStudy (cwrw238):

you should get 6.389 for that

OpenStudy (anonymous):

oooh ok i see what i did before, before i was using ln() to get rid of the e and to bring the exponent of 2 down, i thought thats what i had to do because of some stuff i was doing earlier that required that method (think derivatives)

OpenStudy (cwrw238):

just plug t = 2 into the formula

OpenStudy (anonymous):

ok so for t=2 you get 6.389 t=4 you get 53.598

OpenStudy (anonymous):

so next is \[\frac{ 53.589-6.389 }{ 2 }\]

OpenStudy (anonymous):

giving me 23.6045

OpenStudy (anonymous):

so my avg velocity is 23.6 um/s

OpenStudy (cwrw238):

i got 23.796

OpenStudy (anonymous):

hm let me try again

OpenStudy (cwrw238):

no sorry you are right

OpenStudy (anonymous):

hm not sure how

OpenStudy (anonymous):

oh haha alright :D

OpenStudy (cwrw238):

23.6 is correct

OpenStudy (anonymous):

awesome so that is my avg velocity right?

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

oh wow, i was so close the first time, i just added a weird step that ruined the rest

OpenStudy (cwrw238):

the units are um per second or um s(-1

OpenStudy (cwrw238):

or as you wrote it um/s

OpenStudy (anonymous):

yeah i got the unit part, i just didnt write it that last time

OpenStudy (cwrw238):

ok - average velocity = total distance / time taken

OpenStudy (cwrw238):

gotta go - good luck

OpenStudy (anonymous):

alright thanks for the help!